After graduation, you work as a Financial Analyst in a reputable multinational company in Kuala Lumpur. Your superior has assigned you a new task. You are required to analyse the following bonds.

Company | Petron Berhad | Shell Corporation

Coupon rate per annum | 10% | 10%

Maturity in years | 15 | 10

Face value per bond | $1,000 | $1,000

Yield to maturity per annum | 11% | 13%

Current selling price per bond | $900 | $1,100

Based on the above information, you are required to:

a) Compute the current bond prices for both companies if the interest payment is once a year.

b) Based on the above result, provide your conclusion.

c) Discuss three main differences between conventional bond and Islamic bond


1
Expert's answer
2021-07-06T12:06:23-0400

Bond price for Petrol Berhad

Annual interest payment = "0.1 \u00d7 1000= 100"

Price of bond;

="\frac{100}{1.11}+ \frac{100}{1.11^2}+\frac{100}{1.11^3}.....+\frac{100+1000}{1.11^{15}}"

= "928.0913"


Bond Price for Shell Corporation

Annual Coupon interest="0.1\u00d7 1000"

="100"

Bond Price = "\frac{100}{ 1.13} + \frac{100}{1.13^2}+\frac{100}{1.13^3}..+ \frac{100+1000}{1.13^{10}}"

="837.2127"


The bond by Petrol Berhad is a better one because it has a higher bond price value at the end of the maturation period compared to the bond by Shell Corporation.It is also less riskier compared to the one by Shell Corporation.


c)Differences between conventional and Islamic bonds

  • The face value of Islamic bond is based on the market value of the underlying asset, whereas the face value of a bond is based on the issuer’s credit-worthiness.
  • Islamic bond is governed by the Sharia law such that it can't deal in assets such as gambling, alcohol or weapons whereas conventional bonds finance any asset that comply with the local legislation.
  • Islamic bonds are backed by tangible assets and investors are given ownership of the assets whereas conventional bonds are debt obligations.